# The WINTER Observatory: A One-Degree InGaAs Survey Camera to study the Transient Infrared Sky

### v1

@@ VI.3 Data reduction pipeline
The WINTER data reduction and transient alert pipeline build upon the work for the ZTF and Palomar Gattini IR projects. The data reduction pipeline, named the Modular Image Reduction and Analysis Resource777https://github.com/winter-telescope/mirar (mirar) is a open-source, modular, Python 3.11-based pipeline (V. Karambelkar and R.D. Stein et al., in prep). mirar processes both SUMMER and WINTER images, going through a series of standard infrared data reduction steps:
@@
Calibration: Dark subtraction, flat-field correction, removal of sensor pattern noise through Fourier filtering, and subtraction of the modeled sky background with the SExtractor package (Bertin1996).
@@
Astrometric calibration: Each image is astrometrically calibrated using astrometry.net (astrometry.net) and a cross-matched catalogue of sources detected in Gaia (Gaia) and 2MASS (Cutri2003). Bad images are once again rejected at this step.
@@
Stacking: Each WINTER image is split into a set of shorter exposures dithered around a central point, to enable long exposure times with the bright infrared sky background and reconstruction of the point spread function (PSF). The dithered exposures are stacked using the Swarp algorithm (swarp).
@@
Difference imaging: Nightly images are subtracted with the ZOGY package (zogy) from UKIRT (J and Hs), Pan-STARRS (Y), or WINTER (Y, J, and Hs) reference images, where available.
@@
Sources classified as “real” are uploaded with the Apache Avro888https://avro.apache.org/ data serialization system to IPAC at Caltech, where they are sent to the Kowalski alert broker and the Skyportal frontend (skyportal) for alert visualization and follow-up coordination (Figure 17). Candidates appear daily on the Fritz web portal, which is a specific instance of Skyportal with a Kowalski backend, for human vetting and coordination of follow-up observations.
@@ VII.1 Photometric depth
The original design for WINTER predicted a limiting magnitude of for a 5-minute exposure under ideal observing conditions (Frostig:2020). Multiple factors in the realization of the instrument led to the limiting magnitudes of WINTER to be in a 16-minute exposure (Figure 18). The most significant change is the effective QE, reduced from the designed QE (see Section III.4.2). Additionally, increased dark current, higher gain, and a larger focused spot size compared to the original design introduce secondary performance degradations (Table 1). There are also significant variations in performance between the six sensors, caused by differences in quantum efficiency, gain, dark current at the final operating temperatures, and hybridization issues present in some sensors (e.g., the glow spots seen in Figure 3). This leads to a median depth in a 16-minute exposure of 17.6-18.6 in (8 dithers), 17.6-18.7 in (8 dithers), and 16.7-17.7 in (15 dithers) for the six sensors for MJD 60849 - 60856. The data reduction pipeline is optimized around -band imaging, with and optimization planned. The depths also vary significantly with the conditions of the sky background.
@@ VII.2 Early science with WINTER
WINTER observations balance a mix of long-term surveys with ToO interrupts (Figure 16 shows the current history of WINTER observations). The core WINTER community surveys target regions of the sky predicted to yield the most infrared transients and variables in a blind search. This includes a 14-day cadence Galactic plane survey in - and -bands and a series of galaxy-targeted surveys. Galaxies within 25 Mpc, ULIRGS, and galaxy clusters are surveyed at a 10-day cadence in -band. For the first year of operations, very nearby galaxies, such as M31, were monitored daily in -, -, or -bands. In addition to the core surveys, WINTER builds up -band reference images of the sky, prioritizing fields above 60∘ declination where there are no UKIRT or VISTA images. The surveys are regularly interrupted by ToO requests with minutes-scale response times. The WINTER collaboration has published over seventy General Coordinates Network (GCN) circulars, including regular -band detections of gamma-ray bursts (e.g., 2025GCN.42562....1M) and detections of potential counterparts to neutrino alerts (e.g., 2024GCN.35584....1S).
WINTER sets itself apart from other infrared surveys with its time-domain infrastructure, enabling rapid alerts and prompt spectroscopic follow-up of transient sources. For example, WINTER observed a young stellar object outburst (Figure 19). The team followed up the event with the MMIRS spectrograph on the MMT (MMIRS; MMIRS_drp) and classified the event as an FU Ori type outburst following the classification scheme in Connelley:2018. This event is also observed and studied in (Gaia18dvy). WINTER also observed a highly reddened galactic classical nova (2024ATel16658....1W) (Figure 19). Follow-up spectroscopy with the FIRE spectrograph on the Magellan telescopes (FIRE) shows the presence of carbon lines, which could indicate the event is a Fe II classical nova (NIRnovae). Other science with WINTER includes infrared coverage of supernovae, which are well covered in optical surveys but not infrared surveys (Figure 20), and luminous red novae in nearby galaxies.
The collaboration has released three studies with early WINTER data. Frostig:2025 present NIR follow-up of the GW candidate S250206dm, covering 43% of the 50% localization region but finding no viable kilonova candidates, consistent with the large inferred distance of 373 Mpc. Karambelkar:2025 report the slow eruption of an early-AGB star in M31 (WNTR23bzdiq/WTP19aalzlk), which exhibited a decade-long optical/IR brightening and spectroscopic properties resembling stellar merger transients, possibly signaling the onset of common-envelope evolution. Finally, Frostig:2025b characterize two infrared FU Ori outbursts discovered by WINTER and NEOWISE, including one new confirmation of an embedded Class I protostar, demonstrating WINTER’s ability to uncover obscured protostellar eruptions inaccessible to optical surveys.
@@ VIII Conclusions
WINTER is running nightly robotic near-IR surveys since June 2023. Through a mix of extragalactic and Galactic plane surveys, targeted follow-up, and tiled searches, the project is enabling studies of reddened and intrinsically red transients and variables. The time-domain infrastructure supports rapid follow-up of new discoveries. Recent highlights from early WINTER science include searches for kilonovae from gravitational-wave triggers, evidence of common envelope evolution in AGB stars, and characterization of outbursts and mass accretion in young stellar objects.
While WINTER’s sensors did not meet design specifications, the project demonstrates the promise and need for wide-format InGaAs sensors at an accessible cost. Extensive testing presented here found decreased quantum efficiency from the design specification (from 80% to 10%) due to an issue in hybridization or processing of this lot of sensors. The InGaAs technology has no fundamental flaws, and the read out integrated circuit, and custom read out electronics all performed near expectations. The decreased quantum efficiency, along with secondary effects from increased dark current, gain, and focus spot size, reduced the on-sky depth by 2 magnitudes. However, a broad range of science is underway at the present average depth of mag, which is significantly fainter than historical IR time domain surveys.
WINTER studies complement and prepare for other infrared surveys. Future surveys such as the Roman Space Telescope and Cryoscope (Cryoscope) will enable wider and deeper near-IR observations, extending studies of these transients to greater depths. In the meantime, WINTER is developing near-IR infrastructure, surveys, and science for the time-domain era.
@@ IX Acknowledgments
This work made use of Astropy:999http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (astropy:2013; astropy:2018; astropy:2022).

### v2

@@ VI.3 Data reduction pipeline
The WINTER data reduction and transient alert pipeline build upon the work for the ZTF and Palomar Gattini IR projects. The data reduction pipeline, named the Modular Image Reduction and Analysis Resource777https://github.com/winter-telescope/mirar (mirar) is a open-source, modular, Python 3.11-based pipeline (V. R. Karambelkar and R. D. Stein et al., in prep). mirar processes both SUMMER and WINTER images, going through a series of standard infrared data reduction steps:
@@
Calibration: Dark subtraction, flat-field correction, removal of sensor pattern noise through Fourier filtering, and subtraction of the modeled sky background with the SExtractor package (Bertin and Arnouts, 1996).
@@
Stacking: Each WINTER image is split into a set of shorter exposures dithered around a central point, to enable long exposure times with the bright infrared sky background and reconstruction of the point spread function (PSF). The dithered exposures are stacked using the Swarp algorithm (Bertin et al., 2002).
@@
Difference imaging: Nightly images are subtracted with the ZOGY package (Vreeswijk and Paterson, 2021) from UKIRT (J and Hs), Pan-STARRS (Y), or WINTER (Y, J, and Hs) reference images, where available.
@@
Sources classified as “real” are uploaded with the Apache Avro888https://avro.apache.org/ data serialization system to IPAC at Caltech, where they are sent to the Kowalski alert broker and the Skyportal frontend (van der Walt et al., 2019) for alert visualization and follow-up coordination (Figure 17). Candidates appear daily on the Fritz web portal, which is a specific instance of Skyportal with a Kowalski backend, for human vetting and coordination of follow-up observations.
@@ VII.1 Photometric depth
The original design for WINTER predicted a limiting magnitude of for a 5-minute exposure under ideal observing conditions (Frostig et al., 2020). Multiple factors in the realization of the instrument led to the limiting magnitudes of WINTER to be in a 16-minute exposure. The most significant change is the effective QE, reduced from the designed QE (see Section III.4.2). Additionally, increased dark current, higher gain, and a larger focused spot size compared to the original design introduce secondary performance degradations (Table 1). There are also significant variations in performance between the six sensors, caused by differences in quantum efficiency, gain, dark current at the final operating temperatures, and hybridization issues present in some sensors (e.g., the glow spots seen in Figure 3). This leads to a median depth in a 16-minute exposure of 17.6-18.6 in (8 dithers), 17.6-18.7 in (8 dithers), and 16.7-17.7 in (15 dithers) for the six sensors for MJD 60849 - 60856 (Figure 18). The data reduction pipeline is optimized around -band imaging, with and optimization planned. The depths also vary significantly with the conditions of the sky background.
@@ VII.2 Early science with WINTER
WINTER observations balance a mix of long-term surveys with ToO interrupts (Figure 16 shows the current history of WINTER observations). The core WINTER community surveys target regions of the sky predicted to yield the most infrared transients and variables in a blind search. This includes a 14-day cadence Galactic plane survey in - and -bands and a series of galaxy-targeted surveys. Galaxies within 25 Mpc, ULIRGS, and galaxy clusters are surveyed at a 10-day cadence in -band. For the first year of operations, very nearby galaxies, such as M31, were monitored daily in -, -, or -bands. In addition to the core surveys, WINTER builds up -band reference images of the sky, prioritizing fields above 60∘ declination where there are no UKIRT or VISTA images. The surveys are regularly interrupted by ToO requests with minutes-scale response times. The WINTER collaboration has published over seventy General Coordinates Network (GCN) circulars, including regular -band detections of gamma-ray bursts (e.g., Mo et al. 2025) and detections of potential counterparts to neutrino alerts (e.g., Stein et al. 2024).
WINTER sets itself apart from other infrared surveys with its time-domain infrastructure, enabling rapid alerts and prompt spectroscopic follow-up of transient sources. For example, WINTER observed a young stellar object outburst (Figure 19). The team followed up the event with the MMIRS spectrograph on the MMT (McLeod et al., 2012; Chilingarian et al., 2015) and classified the event as an FU Ori type outburst following the classification scheme in Connelley and Reipurth (2018). This event is also observed and studied in (Szegedi-Elek et al., 2020). WINTER also observed a highly reddened galactic classical nova (Wang and Bellm, 2024) (Figure 19). Follow-up spectroscopy with the FIRE spectrograph on the Magellan telescopes (Simcoe et al., 2013) shows the presence of carbon lines, which could indicate the event is a Fe II classical nova (Banerjee and Ashok, 2012). Other science with WINTER includes infrared coverage of supernovae, which are well covered in optical surveys but not infrared surveys (Figure 20), and luminous red novae in nearby galaxies.
The collaboration has released three studies with early WINTER data. Frostig et al. (2025b) present near-IR follow-up of the GW candidate S250206dm, covering 43% of the 50% localization region but finding no viable kilonova candidates, consistent with the large inferred distance of 373 Mpc. Karambelkar et al. (2025) report the slow eruption of an early-AGB star in M31 (WNTR23bzdiq/WTP19aalzlk), which exhibited a decade-long optical/IR brightening and spectroscopic properties resembling stellar merger transients, possibly signaling the onset of common-envelope evolution. Finally, Frostig et al. (2025a) characterizes two infrared FU Ori outbursts discovered by WINTER and NEOWISE, including one new confirmation of an embedded Class I protostar, demonstrating WINTER’s ability to uncover obscured protostellar eruptions inaccessible to optical surveys.
@@ VIII Conclusions
WINTER has been running nightly robotic near-IR surveys since June 2023. Through a mix of extragalactic and Galactic plane surveys, targeted follow-up, and tiled searches, the project is enabling studies of reddened and intrinsically red transients and variables. The project’s time-domain infrastructure supports rapid follow-up of new discoveries. Recent highlights from early WINTER science include searches for kilonovae from gravitational-wave triggers, evidence of common envelope evolution in AGB stars, and characterization of outbursts and mass accretion in young stellar objects.
While WINTER’s sensors did not meet design specifications, the project demonstrates the promise and need for wide-format InGaAs sensors at an accessible cost. Extensive testing presented here found decreased quantum efficiency from the design specification (from 80% to 10%) due to an issue in hybridization or processing of this lot of sensors. The InGaAs technology has no fundamental shortcomings, and the read out integrated circuit, and custom read out electronics all performed near expectations. The decreased quantum efficiency, along with secondary effects from increased dark current, gain, and focus spot size, reduced the on-sky depth by 2 magnitudes. However, a broad range of science is underway at the present average depth of mag, which is significantly fainter than historical IR time-domain surveys.
WINTER studies complement and prepare for other infrared surveys. Future surveys such as the Roman Space Telescope and Cryoscope will enable wider and deeper near-IR observations, extending studies of these transients to greater depths. In the meantime, WINTER is developing near-IR infrastructure, surveys, and science for the time-domain era.